# Towards Accurate and Interpretable Neuroblastoma Diagnosis via Contrastive Multi-scale Pathological Image Analysis

### v1

@@ 4.6.1 Model Performance Ablation
Our ablation experiments further analyzed the impact of the KAN network, CDFA component, and CMSA on classification performance across both datasets. We assessed the performance of the CMSwinKAN model after restoring the KAN network, replacing the CDFA component with a linear layer, removing the CMSA module, and comparing it to the complete CMSwinKAN model. Table 8 confirms that these three modules positively influence the model’s classification performance. Specifically:
Impact of the KAN Network: In the SwinKANsformer Block, the KAN network replaces the traditional MLP layer, enhancing the model’s ability to extract nonlinear features and improving interpretability. The conventional linear layer classification head is replaced with a more efficient KAN network. If the KAN network is restored and the MLP layer and linear classification head are used instead, the classification performance significantly declines. Specifically, ACC drops from 93.58% to 92.29%, BACC from 93.42% to 92.09%, and KAPPA and F1 scores decrease from 92.12% and 93.57% to 90.93% and 92.28%, respectively. This decline is due to a significant reduction in feature expression capability when reverting to the MLP layer and linear classification head, which fails to adequately extract fine-grained features in pathological slice images, leading to overly similar learned features that easily confuse different cell types.
@@
Replacing the CDFA Module with a Linear Layer: The CDFA module enhances the model’s learning ability for information from global to microscopic scales through feature comparison across different scales. When the CDFA module is replaced with a linear layer, effective interaction between different scales is lost, resulting in significant feature information loss and a marked decline in all metrics. Specifically, KAPPA decreases by 1.29%, and F1 scores and other metrics drop by 0.7%. This highlights the crucial role of the CDFA module in multi-scale feature fusion, significantly aiding the model’s ability to integrate high-level semantic information and detailed edge information in pathological images.
@@
Table 8: Ablation studies of CMSwinKAN on our private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC(%)
w/o KAN 92.29 92.09 90.93 92.28 99.39
w/o CDFA 92.89 92.73 90.93 92.88 99.44
w/o CMSA 92.50 92.24 90.85 92.49 99.42
CMSwinKAN 93.58 93.42 92.12 93.57 99.53
Table 9: Ablation studies of CMSwinKAN on the public BreakHis dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC(%)
w/o KAN 80.66 79.59 77.31 80.90 97.81
w/o CDFA 82.18 82.41 75.75 82.60 98.14
w/o CMSA 79.64 81.40 77.44 80.06 97.78
CMSwinKAN 87.79 87.56 85.69 88.04 98.80
Removing the CMSA Module: The CMSA module combines the CDFA module with a learnable weight allocator, adjusting the contribution distribution of multi-scale features to capture long-range dependencies and local features in pathological slice images while avoiding contamination from scales with excessive low-density information. When the CMSA module is removed, the complete loss of multi-scale details leads to a significant decline in the model’s ability to capture global and local features, resulting in noticeable reductions in performance metrics. Specifically, ACC decreased by 1.15%, BACC by 1.28%, and KAPPA by 1.38%.
Notably, the performance decline from removing the CMSA module is more pronounced than replacing the CDFA module with a linear layer, reflecting the effectiveness of the CMSA module. The multi-scale contribution weight allocation significantly supports the model in extracting effective high-density feature information, thereby enhancing the powerful capabilities of multi-scale feature learning when combined with the CDFA module.
4.6.2 Validation of Model Generalization Ability
Additionally, we conducted the same ablation experiments on the publicly available breast tissue histopathological image classification dataset, BreakHis, which features an imbalanced class distribution, further reflecting the impact of different components on model classification performance. These experiments further validate the significant influence of each key component on classification accuracy. As shown in Table 9, removing any component leads to a marked decline in classification metrics, highlighting the importance of these components in feature extraction and expression capabilities, as well as the necessity of effectively utilizing high-level semantics and detailed edges in pathological images to simulate clinical pathology diagnostics, thus enhancing the model’s classification ability.
@@ 5 Conclusion
In this paper, we proposed a model named CMSwinKAN for the patch classification task of NB pathological slices stained with H&E. This model addressed several issues present in existing methods, such as insufficient integration of global features and local edge details, poor model interpretability, and the susceptibility of multi-scale features at different stages to contamination from low-density information. Additionally, we designed a soft voting mechanism based on prior knowledge of tissue composition to facilitate an effective transition from patch-level classification to WSI-level classification.
We evaluated the performance of the proposed CMSwinKAN model against other classical and advanced methods for both patch-level and WSI-level classification on the private PpNTs dataset and the publicly available BreakHis dataset. The experimental results demonstrated that CMSwinKAN significantly outperforms existing pathological image classification methods on these datasets and markedly enhances model performance in WSI classification tasks through the soft voting mechanism, achieving excellent results across nearly all evaluation metrics. Furthermore, we conducted ablation experiments to validate the importance of key components within the model, and the results confirmed that these components contributed positively and significantly to model performance.

### v2

@@ 4.5 Ablation Experiments
Our ablation experiments further analyzed the impact of the KAN network, CDFA component, and CMSA on classification performance across both datasets. Table 6 confirmed that these three modules positively influence the model’s classification performance. Specifically:
Impact of the KAN Network:In the SwinKANsformer Block, the KAN network replaces the traditional MLP layer, improving nonlinear feature extraction and interpretability. Switching to the MLP layer leads to a drop in classification performance: ACC decreases from 93.58% to 92.29%, BACC from 93.42% to 92.09%, and KAPPA and F1 scores decline from 92.12% and 93.57% to 90.93% and 92.28%. This drop occurs due to a reduced ability to capture fine details in pathological images, leading to cell type confusion.
@@
Replacing the CDFA Module with a Linear Layer:The CDFA module improves the model’s learning by effectively comparing features across various scales. Replacing it with a linear layer leads to significant feature loss, with KAPPA decreasing by 1.29% and F1 scores dropping by 0.7%. This underlines the CDFA module’s essential role in multi-scale feature fusion, enhancing the model’s integration of high-level semantic and detailed edge information in pathological images.
Removing the CMSA Module: The CMSA module enhances the CDFA module by optimizing multi-scale feature contributions and capturing both long-range dependencies and local features in pathological slice images while reducing contamination from low-density scales. Removing the CMSA module significantly decreases performance metrics: BACC by 1.28%, and KAPPA by 1.38%. This decline is greater than that seen when the CDFA module is replaced with a linear layer, underscoring the CMSA module’s crucial role in high-density feature extraction and multi-scale learning.
@@
Table 6: Ablation studies of CMSwinKAN on our private PpNTs and public BreakHis datasets.
Dataset PpNTs BreakHis
Model ACC BACC KAPPA F1 AUROC ACC BACC KAPPA F1 AUROC
w/o KAN 92.29 92.09 90.93 92.28 99.39 80.66 79.59 77.31 80.90 97.81
w/o CDFA 92.89 92.73 90.93 92.88 99.44 82.18 82.41 75.75 82.60 98.14
w/o CMSA 92.50 92.24 90.85 92.49 99.42 79.64 81.40 77.44 80.06 97.78
CMSwinKAN 93.58 93.42 92.12 93.57 99.53 87.79 87.56 85.69 88.04 98.80
@@ 5 Conclusion
This paper proposed a model named CMSwinKAN for the pathological-grade diagnosis NB pathological slices stained with H&E. The model innovatively integrates multi-scale features and comparative learning strategies, effectively addressing the key issues such as insufficient feature integration of existing methods, poor model interpretability, and the easy contamination of multi-scale features by low-density information. Additionally, a heuristic soft voting mechanism was designed, guided by clinical insights, to achieve an efficient transition from patch-level classification to WSI-level classification. Experiments on private PpNTs and public BreakHis datasets demonstrate that this method significantly outperforms the prior art. Particularly, the interpretable output of the model is highly consistent with the pathological diagnostic criteria, providing reliable support for clinical AI-assisted diagnosis. Future work will focus on multi-center data verification and exploring the fusion application with molecular diagnosis to further enhance the clinical practical value.